Isabel drove 737 miles in 11 hours.At the same rate,how long would it take her to drive 469 miles?

Answer by Love math(28)

(737/11)=(469/h)
cros mutipfly
(737*h)=(11*469)
737h = 5.159
h = 5.159/737
7 hours